zarkam21:

@nnesha my teacher said some of these are wrong can you help me correct them (4 questions)

zarkam21:

@Nnesha

zarkam21:

1 attachment
zarkam21:

Could you tell me why I did this wrong

zarkam21:

@Nnesha

Nnesha:

hmm it's larger for sure..

zarkam21:

maybe its choice b then instead of C maybe?

Nnesha:

yeah but i don't understand why it's not the C since it's negative the image is behind the mirror

Nnesha:

or maybe image distance is not negative

zarkam21:

1 attachment
zarkam21:

this is what it says in my lesson

Nnesha:

\[\rm -6.1=\frac{d_i}{d_o}\] if d_o is always positive (according to the table) then for this equation image distance needs to be negative

zarkam21:

meaning that the answer is B?

zarkam21:

1 attachment
zarkam21:

I hvae to fill out this table

Nnesha:

no i was trying to figure out why it is not C. BUT if is not C then ofc it is definitely B. other 2 are for smaller

zarkam21:

SO would this be the right answer for why it is wrong?

Nnesha:

According to this website http://physics.bu.edu/py106/notes/Spherical.html `When the image distance is positive, the image is on the same side of the mirror as the object, and it is real and inverted. When the image distance is negative, the image is behind the mirror, so the image is virtual and upright. A negative m means that the image is inverted. Positive means an upright image.`

zarkam21:

My answer was wrong because I incorrectly came to the conslusion that imageis behind the mirror. The correct answer is that it is large and in front.

zarkam21:

I need a reason as to why it is the correct answer :(

zarkam21:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @zarkam21 My answer was wrong because I incorrectly came to the conslusion that imageis behind the mirror. The correct answer is that it is large and in front. \(\color{#0cbb34}{\text{End of Quote}}\)

zarkam21:

you think this is good enough

Nnesha:

no

zarkam21:

1 attachment
Nnesha:

If it was me I will put why i Picked the option C in first place. and will ask teacher why it is not C.

zarkam21:

this is also part of my lesson

zarkam21:

so from this because it is negative wouldnt it be D , because a negative means the image is small and behind the mirror

Nnesha:

oh okay one second

Nnesha:

no sign doesn't explain if the image is going to be smaller or larger the number itself will rperesent that

Nnesha:

we know that image is 100% larger because 6.1>1

zarkam21:

Here is an example of a negative image distance

zarkam21:

1 attachment
zarkam21:

maybe this will help?

zarkam21:

this is about negative mag.

zarkam21:

1 attachment
Nnesha:

\[\large\rm m=-\frac{d_i}{d_o} \rightarrow -m=\frac{d_i}{d_o}\] if we move the negative sign then it makes sense that d_i is positive which means the image is in front of the mirror. However if we look at this equation mathematically if d_o needs to be positive then di must be negative. so I will ask teacher

zarkam21:

okay

zarkam21:

I will ask the teacher for this one

Nnesha:

\[\rm d_o= \frac{-m}{d_i}\] in order to cancel out negative sign di should be negative

Nnesha:

or maybe we just have to ignore that

zarkam21:

well they are saying I am wrong and I, myself do not understand why . So if you can figure out why I would sppreciate it. If you dont understand what I did wrong either, I will ask the teacher

Nnesha:

B is also wrong ??

zarkam21:

No no C is

zarkam21:

so b is the right answer ?

Nnesha:

if it marked C wrong then yeah B needs to be the answer....

zarkam21:

OKay sounds good. can we move on to the next one. (There are 4 )

Nnesha:

will see. the last time i did these types of question were like years+infinity ago.. but will try

zarkam21:

1 attachment
Nnesha:

object and image are on the other sides so image distance needs to be negative

Nnesha:

opposite sides*

zarkam21:

SO prob choice D

Nnesha:

what does the red cross mark represent ? i'm confused

zarkam21:

the answer is wrong

zarkam21:

that is the answer i submitted but it is wrong

Nnesha:

that's what i thought but after your comment `so prop D` i got confusd for a second \[\rm m=-\frac{d_i}{d_o}= -\frac{-10 cm}{20cm}\]

zarkam21:

-0.5

zarkam21:

but i got that wrong

Nnesha:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Nnesha that's what i thought but after your comment `so prop D` i got confusd for a second \[\rm m=-\frac{d_i}{d_o}= -\frac{-10 cm}{20cm}\] \(\color{#0cbb34}{\text{End of Quote}}\) \[\rm m=-\frac{d_i}{d_o}= \color{blue}{-}(\frac{\color{ReD}{-}10 cm}{20cm})\]

zarkam21:

ohhhh its 0.5

zarkam21:

because of the double negative

Nnesha:

yes :=))

zarkam21:

yay

zarkam21:

would this be a good explanation for why I got it wrong? My answer was wrong because my calculations were wrong. m=− di/do=−(-10cm)/(20cm). The answer is 0.50 because of the double negative and I put -0.50

Nnesha:

yes you can say that but my explanation will be different

Nnesha:

is this a online class ?

zarkam21:

yes it is

zarkam21:

what would your explanation be

Nnesha:

I don't think your mistake was multiplying two negative values.. You didn't realize that the image distance needs to be negative. so if it was me i will mention that to my teacher that I couldn't figure out that Di needs to be negative because image and the object are on opposite side that's how your teacher will understand that you're trying your best but having difficult time to catch the key points maybe she/he will help you with that

zarkam21:

Yeah true

zarkam21:

shall we continue to the last one

Nnesha:

last one ? yayya sure

zarkam21:

actually second to last one

zarkam21:

there is one more after this

zarkam21:

1 attachment
Nnesha:

hmm what do you think the answer should be now after figuring out the first question we did together ?

zarkam21:

D

Nnesha:

1526307318-5af999d93bef0945d7824c00-image.png

zarkam21:

D?

Nnesha:

well first of all m >1 or m <1 ??

zarkam21:

Yes that is true

zarkam21:

it is less than

Nnesha:

hmm ignore the sign for now

Nnesha:

from yesterday post ` when m is negative => the image is inverted` `when m is positive => the image is upright` `when m is greater than 1 => image is larger than the object` `when m is less than 1 => image is smaller than the object``

zarkam21:

when m is less than 1 => image is smaller than the object`

zarkam21:

thsi is the one we should look at because it is smaller than 1

Nnesha:

If i'm not wrong we need to look at sign and the value separately because they both represent different things

zarkam21:

inverted and smaller than

zarkam21:

when m is negative => the image is inverted when m is less than 1 => image is smaller than the object`

Nnesha:

m= 2 ignoring the sign is 2<1 or 2>1 ?

zarkam21:

2>1

Nnesha:

yes so m is greater than m

zarkam21:

okay so fromt hat we can see that it is larger than the object

Nnesha:

yes

zarkam21:

so choice c

zarkam21:

we have one more

zarkam21:

1 attachment
Nnesha:

yes C because m is negative and di is positive so it should be real image

Nnesha:

ohh we did that one yesterday hmm I was confused about that one bec if we just look at the table both virtual are negative..

zarkam21:

maybe retrying witht he lessons I showed ? :/ (

Nnesha:

where are the lessons ?

zarkam21:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @zarkam21 \(\color{#0cbb34}{\text{End of Quote}}\)

zarkam21:

hold on

zarkam21:

1 attachment
zarkam21:

1 attachment
zarkam21:

1 attachment
zarkam21:

1 attachment
zarkam21:

1 attachment
zarkam21:

1 attachment
Nnesha:

hmm I agreed to D after watching this video yesterday https://www.khanacademy.org/science/physics/geometric-optics/lenses/v/thin-lens-equation-and-problem-solving @8.25

zarkam21:

Did you find the answer

zarkam21:

maybe A

zarkam21:

because that would be the opposite

Nnesha:

no i won't pick A

zarkam21:

Then?

Nnesha:

https://prnt.sc/jicf38

Nnesha:

the negative indicates it is virtual

zarkam21:

SO B

Nnesha:

i'm still looking for more info maybe 5 minutes ?

zarkam21:

yeah sure no prob take your time

Nnesha:

if the image distance is negative the image will be virtual according to that ^^both can have virtual image